OpenStudy (zubhanwc3):

a couple of calculus questions regarding velocity, and position, acceleration type of stuff

OpenStudy (zubhanwc3):

1. A particle is moving along the x-axis so that at time t its acceleration is . At time t = ½, the velocity, v of the particle is ½. a) Find the velocity of the particle at any time t. b) Find the minimum velocity of the particle. c) Find the equation for the position x(t) if x(0) = 0. d) What is the first time t > 0 that the particle returns to the origin? 2. A particle is moving with velocity: with distance, s measured in meters, left or right of zero, and t measured in seconds, with t between 0 and 8 seconds inclusive. The position at time t = 0 sec is 1 meter right of zero. Find each of the following. Give units in the answer. a) The initial acceleration b) The average velocity over the interval 0 to 8 seconds c) The instantaneous velocity at time 5 secs d) The time(s) when the particle is at rest e) The time interval(s) when the particle is moving left f) The time interval(s) when the particle is moving right g) The speed of the particle at time 4 secs h) The time interval(s) when the particle is (i) going faster (ii) slowing down i) Find the total distance the particle has traveled between 0 and 8 seconds

OpenStudy (nincompoop):

what is the rate of acceleration?

OpenStudy (zubhanwc3):

i dont know, this is all the info i have.

zepdrix (zepdrix):

1. A particle is moving along the x-axis so that `at time t its acceleration is ` . At time t = ½, the velocity, v of the particle is ½. Are you sure? it looks like something didn't paste correctly.

OpenStudy (zubhanwc3):

ah your right. a(t) = picospi

OpenStudy (nincompoop):

1/2 velocity is incorrect description velocity is a vector so it needs to have both magnitude and direction

OpenStudy (zubhanwc3):

i dunno why it didnt copy tho.

OpenStudy (nincompoop):

you're not allowed to copy and paste your homework here btw

OpenStudy (zubhanwc3):

its my review questions for my test.

OpenStudy (zubhanwc3):

and im confused.

OpenStudy (zubhanwc3):

alright, i dont mind that

OpenStudy (zubhanwc3):

but if you want the file, i can attach it. for proof

zepdrix (zepdrix):

So the acceleration is given to be constant. Have you been introduced to integration yet? Or just the idea of anti-differentiation?

OpenStudy (zubhanwc3):

yes, i have

zepdrix (zepdrix):

\[\Large a(t)\quad=\quad \pi \cos \pi\]Taking the anti-derivative of acceleration should give us the velocity function. Do you understand why that is? Do you know how to find the anti-derivative of this function? :) *cough* anti-derivative of a constant? *cough*

OpenStudy (zubhanwc3):

because the derivative of velocity would equal the acceleration? and the anti-derivative would be -pix

zepdrix (zepdrix):

Well we wouldn't be getting an x, we'd get a t in this case. But yes! You have the right idea. We'd also want a constant of integration.\[\Large\bf\sf v(t)\quad=\quad -\pi t+c\]

OpenStudy (zubhanwc3):

ah right, forgot about that:P

zepdrix (zepdrix):

We can use our initial data to solve for the unknown constant.\[\Large\bf\sf v\left(\frac{1}{2}\right)\quad=\quad \frac{1}{2}\]We'll plug this into our velocity function to solve for c.

OpenStudy (zubhanwc3):

so 1/2 = pi(1/2) + c so c= (pi)(1/2) - 1/2)

OpenStudy (zubhanwc3):

which is equal to 1.0708

zepdrix (zepdrix):

we had a negative in front of our pi didn't we? 1/2 = -pi(1/2) + c so c= pi(1/2) + 1/2 yes? :o

OpenStudy (zubhanwc3):

ah, so it would be 2.0708 instead

zepdrix (zepdrix):

So for part a, we want to velocity function with our constant plugged in.\[\Large\bf\sf v(t)\quad=\quad -\pi t+\color{orangered}{c}\]

zepdrix (zepdrix):

want the*

zepdrix (zepdrix):

It probably looks nicer if we don't approximate.\[\Large\bf\sf v(t)\quad=\quad -\pi t+\color{orangered}{\frac{1}{2}\pi+\frac{1}{2}}\]But do whatever makes more sense to you.

OpenStudy (zubhanwc3):

so the answer to a is) v(t) = -pi(t) + 2.0708

zepdrix (zepdrix):

ya looks good c:

OpenStudy (zubhanwc3):

how would i fine the minimum velocity, by plugging in 0?

zepdrix (zepdrix):

If the acceleration is `constant` then the velocity is `linear`. Since the velocity function has a negative coefficient in front, it will decrease at a constant rate. The velocity function won't have any critical point which we could call a minimum. If you wanted to be a little more rigorous you could take the derivative of v(t) and then set it equal to zero and solve for t to find critical points.

zepdrix (zepdrix):

\[\Large 0\quad=\quad \pi \cos \pi\]No t to solve for :c No critical points. No minimum.

OpenStudy (zubhanwc3):

could the minimum velocity be said to be -infinity?

zepdrix (zepdrix):

Ummmm no I guess we wouldn't want to do that. A minimum should be an actual value. I think... -_-

OpenStudy (zubhanwc3):

so how would i find the position formula

zepdrix (zepdrix):

Do you remember how position relates to velocity? :) *hint hint* similarly to the way velocity relates to acceleration!

OpenStudy (zubhanwc3):

i should take the anti-derivative of v(t)? so that would give me (-pix^2)/2 + c

OpenStudy (zubhanwc3):

since s(0) is 0, c is 0, right?

zepdrix (zepdrix):

Hmm what happened to your previous c in v(t)?

zepdrix (zepdrix):

Let's call our new constant d or something, to avoid confusion.

OpenStudy (zubhanwc3):

er wait, i just took the antiderivative of the original >.>

zepdrix (zepdrix):

\[\Large\bf\sf v(t)\quad=\quad -\pi t+\left(\frac{1}{2}\pi+\frac{1}{2}\right)\]

zepdrix (zepdrix):

So we want to anti-differentiate that.

OpenStudy (zubhanwc3):

so it would be x(1/2pi + 1/2) + x(1/2pix)

OpenStudy (zubhanwc3):

+ c. but c is 0 because of s(0) is equal to 0

zepdrix (zepdrix):

Ok looks like you're on the right track. Where did our negative again though? :( \[\Large\bf\sf s(t)\quad=\quad -\frac{1}{2}\pi t^2+\left(\frac{1}{2}\pi+\frac{1}{2}\right)t+d\]

zepdrix (zepdrix):

s(0)=0 Oh ok good good c=0

OpenStudy (zubhanwc3):

er, its (.5pi + .5) - x(.5pix)

zepdrix (zepdrix):

ok cool so that takes care of part c! :)

OpenStudy (zubhanwc3):

i dont understand part d at all

zepdrix (zepdrix):

So if you look at your position function a sec. What shape does it take? The largest power on x is a square, and it has a negative coefficient. Can you recall what shape it will make? :o

OpenStudy (zubhanwc3):

a bunch of u's. when i graphed the position, it was a linear line, almost like y=x

zepdrix (zepdrix):

Hmm it shouldn't give you that.. It should give you one u. It's a parabola opening downward.

OpenStudy (zubhanwc3):

o, i graphed it wrong:P

zepdrix (zepdrix):

part d is simply asking you to `find the x-intercepts`. They're calling the x-axis the "origin" in this case, which is a little strange. But I guess it makes sense. They want to know when the particle touches back down to the ground or whatever.

zepdrix (zepdrix):

Remember how to find `x-intercepts`? Also called `zeros` or `roots` of a polynomial.

OpenStudy (zubhanwc3):

zeros are 0 and 1.318

zepdrix (zepdrix):

Ok good. and it looks like part d is telling us to ignore one of those points, yes?

zepdrix (zepdrix):

Since one of those points is the starting location.

OpenStudy (zubhanwc3):

so 1.318 is our answer

zepdrix (zepdrix):

At time t = 1.318 seconds, the particle will return to the origin. Ya I think that sounds right! :)

OpenStudy (zubhanwc3):

what about number 2, whats initial acceleration?

zepdrix (zepdrix):

Ahh I dunno. I'm too tired. I need to get some sleep :c Maybe mr poo can help you finish this one up.

OpenStudy (zubhanwc3):

aite, tyvm.

zepdrix (zepdrix):

People hate to read really long posted questions. It might be worth your time to close this thread and open a new question with just #2. You might be able to find some help easier that way c: